Observer EoR ReportBelow is what you submitted to ScienceOps@lbto.org,rgreen@lbto.org on Saturday, January 28, 2012 at 02:55:18

rundates: Jan 26-27, 2012 (first half for both nights)

PrimaryInst: MODS

SecondaryInst:

EoR_program: We obtained multi-slit spectra of young star candidates in two young clusters IC 1795 and W3 south located at the W3 star forming region. Spectra are used to obtain spectral types, to estimate reddening, and to confirm youth of candidate young stars. In order to correctly estimate age and masses of stars we need to know the spectral types and reddening toward each star. We also look for youth indicators, such as H alpha emission from accretion.

EoR_highlight: We obtained good data from second night, and about 50% of targets are bright enough from first night. We think there are interesting stars.

EoR_lossoftime: There were problems due to collimation problem and bad seeing. The secondary seemed to be the source of problem with possibility of dome seeing. Observing staffs here and Tucson tried very hard to identify the source of problem, and tried to improve collimation. During the first night we lost ~1.5 hours during the first half of the night because of the collimation problem. To fix the collimation problem our first target exposures were aborted after 3 exposures (total 7-9 exposures were planned). After lots of tries of finding the source of problem and fixing the secondary problem we tried 2nd field. The second field looked better.

The collimation problem did not go away during the 2nd night, and had to be tweaked after our first target field, but overall 2nd night went a lot better and lost only ~15 minutes during the first half of the night. The MODS alignment script did not correctly calculate position angle shift (wrong sign), so it took 10 image to get it right by manually typing in the PA correction. The problem report was sent to the PI of MODS by Dave Thompson.

EoR_improvements: Better collimation

EoR_comments:

====== TELESCOPE =======

Telescope pointing, tracking, optical performance: Collimation issue was discussed earlier. Guiding could be improved. Observing Support: Excellent.

====== INSTRUMENT ======

Instrument user interface: easy enough to use.

During the night of 20120127, MODS control panel did not show exposure and reading out bars. it did not update the file name either, so we briefly aborted the sequence of taking exposures, refreshed the gui, and continued. The problem did not happen again.

Support Astronomer support: Excellent! Thanks!

Instrument documentation:

Helpful Data reduction software (IRAF, IDL, etc.): quick debias script (python script) worked great. I wish there would be a pipeline in the future, or have a place at wiki where people discuss and post their cookbooks, tips, etc.

===== MISCELLANEOUS =====

Information in advance of observing: It would be helpful to let new observers be advised to prepare observing scripts in advance. I did prepare them before the observing run, but it may not be obvious to some new observers?

Accommodations: Good. More notes: Writing logs in certain suggested format would be a good idea. AZ team writes logs in Twiki site, but new observers may not be aware of it unless the observer has talked with previous observers. Schedule of LBT observing at LBTO site would be helpful. submit_button: Submit Information

S. Kim

-- XiaohuiFan - 29 Jan 2012
Topic revision: r1 - 29 Jan 2012, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback